Tohban Report 2015-04-22

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 15 Apr 2015
End Date: 22 Apr 2015
Tohban: Hazel Bain
Tohban email:
Next Tohban: Nicole Duncan
List all reports


Solar Activity

Solar activity this week was low to moderate. Yesterday (21st Apr) there were 7 M class flares produced by AR NOAA 12322 as it was rotating over the limb. In the last 24 hours solar activity has declined, perhaps due to AR 12322 having rotated over the limb. However the background GOES X-ray flux remains elevated and is currently around the C1 level. There are currently 9 active regions on the disk, the most active of which are NOAA 12324 and NOAA 12325.

How many GOES flares occurred?

 Flares above B, C, M, X class were      2    28     7     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      1    20     5     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     9     2     0

There were RHESSI flares/GOES flares 233 / 37 over the time range 15-Apr-15 22-Apr-15

Memory Management

At the start of the week the SSR was high, reaching 50-60% fill level at the start of the daily pass set. Extra passes, HV commanding and threshold changing (see below) to reduce the noise in D4 and D6 helped reduce the SSR fill level, which is now emptying at the end of the daily pass set.

Spacecraft Status

Data Gaps


Cold Tip Temperature 1 = 108.3K
Cold Tip Temperature 2 = 106.2K
Cold Plate Temperature 1 = 129.9K
Cold Plate Temperature 2 = 128.4K

Detector issues

On the 16th of April, the attenuator logic was changed from detectors 1, 1, 3 and 3 to detectors 1, 1, 7 and 7.

On the 16th of April the high voltage on detector 6 was lowered. The detector remained segmented and both the reset rates and slow valid counts came down. The reset rates started out around 5400 and front slow valids were around 2800 cps.

15-106-08:25:09 /ihvdac detector=6, voltage=153

After lowering by 800 V, the slow valids dropped to around 1800 cps. The G6 high voltage was then lowered by another 20 steps.

15-106-08:27:25 /ihvdac detector=6, voltage=133

After the second move down, the front slow valids dropped to 1200 and the resets dropped to 2000 cps. G6 = 2622 V

On the 16th of April the G4 front slow valids were running close to 1000 cps, so we raised the G4 front slow threshold from 0x50 to 0x60. This brought the counts down to about 600.

15-106-10:04:27 start idib_chg_thrshld(4, FRONT, SLOW, 0x58) 15-106-10:04:38 /IDPUDUMPTABL TABLE=DIBTBL4 15-106-10:04:38 /IDPUTABLE4 OFFSET=FRONTSLOWDAC 15-106-10:04:48 /IDPLOAD VALUE=0x58

15-106-10:05:35 start idib_chg_thrshld(4, FRONT, SLOW, 0x60) 15-106-10:05:42 /IDPUDUMPTABL TABLE=DIBTBL4 15-106-10:05:42 /IDPUTABLE4 OFFSET=FRONTSLOWDAC 15-106-10:05:51 /IDPLOAD VALUE=0x60

Threshold [G4] Front Slow
Initial 0x50
Final 0x60

Other notes

The livetime on detector 8 has decreased around 10% over the last week. The next tohban should keep an eye on the front slow and fast counts.

Spacecraft Management

A yellow "warnings" was added to the the attenuator "in" status to alert operators if the attenuator becomes stuck in state 1. The yellow warning for the CPC temperature was increased from 50 to 55oC (The red level warning remains at 70oC). The yellow warning on solar array 1 (always high due to reflections) was raised to 106oC.

Decimation Active/Vigorous
Night time data (fronts) No nighttime events
Night time data (rears) No nighttime events
Require extra passes? Yes
Requirement for moving pointer? No
Attenuator operation Normal
Personal tools